A restaurant association says the typical household spends a mean of $3568 per year on food away from home. You randomly select 12 households and find out how much each spent on food away from home per year. Can you reject the restaurant association's claim at alpha=0.01? Complete parts a - d.
3678 2205 2444 1917 2295 3456 3900 4267 3004 3157 3233 4510
Write the claim mathematically and identify Ho and Ha. Choose the correct answer
A Ho U = 3422 (claim)
Ha U > 3422
B)Ho > 3422 (Claim)
Ha C ho U > with line under Claim
Ha <
D U = Claim
U =/=
E U =
U < Claim
F U =/=
U = Claim
B) Find the P-value
C Fail to reject or Reject Ho

Answer by stanbon(75887)

A restaurant association says the typical household spends a mean of $3568 per year on food away from home. You randomly select 12 households and find out how much each spent on food away from home per year. Can you reject the restaurant association's claim at alpha=0.01?
---------------------------------------------
Complete parts a - d.
3678 2205 2444 1917 2295 3456 3900 4267 3004 3157 3233 4510
Write the claim mathematically and identify Ho and Ha.
---------------------------------------------
The correct answer
A
Ho: u = 3568 (claim)
Ha: u # 3568
-----
x-bar = 3172
-----
t(3172) = (3172-3568)/[836.52/sqrt(12)] = -1.6399
-------------------------------
B) Find the P-value
p-value = 2*P(-100,-1.6399 when df = 11) = 2*tcdf(-100,-1.6399,11) = 0.1293
-------------------------
C Fail to reject or Reject Ho
Since the p-value is greater than 1% fail to reject Ho.
---------------------
Cheers,
Stan H.
---------------------